# Regularization of the ensemble Kalman filter using a non-parametric, non-stationary spatial model

### v3

@@ 5.5 Effects of covariance regularization
Figure 5(left) shows that the sample (EnKF’s) variances are much noisier than the model variances. This illustrates the fact that the EnKF’s covariance localization cannot improve sample variances, which are quite uncertain for small ensemble sizes. Less noise in the model (LSEF) variances is due to the regularizing effect of using a constrained spatial model. In particular, the local stationarity constraint allows, implicitly, for the use of neighboring covariances in the estimation. The specific mechanism for that is provided by the LSEF’s bandpass filters.
The right panel of Fig. 5 shows that, similar to the left panel, the model correlations contain much less noise than their localized sample counterparts. Another feature seen in this figure is that, in contrast to model correlations, localized sample correlations suffer from a pronounced negative bias in their length scale (see Introduction, item 1). Furthermore, Fig. 5(right) illustrates that fact that the LSEF analysis produces almost noise-free large-distance background-error correlations without any ad-hoc localization (tapering) device. This emphasizes the usefulness of the smooth-local-spectrum constraint (item 4 in section 2.8) and the ability of the neural network to ensure this constraint is met. One can also notice in Fig. 5(right) that the regularizing constraint 1 (section 2.8), which requires the kernel to be locally isotropic, does not lead to isotropic correlations (note the asymmetry in both the true and model correlations).
Averaged over 300 realizations of the truth and the ensemble, the mean absolute error in the model variances was about 1.5 times less than the respective error in the ensemble variances. The mean absolute error in the model spatial correlations was about half of the error in the localized sample correlations. So, both model variances and correlations were substantially more accurate than their localized sample counterparts. This explains the substantial advantage of the LSEF analysis over the EnKF analysis in the experiments presented above in this section.
@@ 5.6 Conclusions on the experiments with static analyses
When the truth was generated by a (parametric) Locally Stationary Convolution Model, the proposed analysis technique performed very well as compared with the competing techniques. The question remains how the proposed approach will perform in a more realistic situation when background errors do not satisfy a Locally Stationary Convolution Model and the true local spectra are not just unavailable but they do not exist. To answer this question we turn to sequential data assimilation.
@@ 6.1 Model of truth and forecast model
${\bf x}_{t}={\bf F}(\boldsymbol{\theta}_{t}){\bf x}_{t-1}+\boldsymbol{\Sigma}_{t}\boldsymbol{\varepsilon}_{t},$ (21)
where ${\bf x}_{t}$ is the spatial field defined on a regular grid on the unit circle (a vector), $t$ is the time index, ${\bf F}(\boldsymbol{\theta}_{t})$ is the forecast model operator, $\boldsymbol{\theta}_{t}$ and $\boldsymbol{\Sigma}_{t}$ are the spatial “parameter” fields defined below in Eqs. (22) and (23), and $\boldsymbol{\varepsilon}_{t}$ is the standard spatio-temporal white noise. Double stochasticity means that both the forcing $\boldsymbol{\varepsilon}_{t}$ and the parameters $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ are random. $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ are postulated to be stationary (in space and time) random fields satisfying the same model Eq. (21) but with constant non-random coefficients $\theta_{\theta},\theta_{\Sigma},\Sigma_{\theta},\Sigma_{\Sigma}$:
$\boldsymbol{\theta}_{t}={\bf F}(\theta_{\theta})\boldsymbol{\theta}_{t-1}+\Sigma_{\theta}\boldsymbol{\varepsilon}_{t}^{\prime},$ (22)
$\boldsymbol{\Sigma}_{t}={\bf F}(\theta_{\Sigma})\boldsymbol{\Sigma}_{t-1}+\Sigma_{\Sigma}\boldsymbol{\varepsilon}_{t}^{\prime\prime},$ (23)
where $\boldsymbol{\varepsilon}_{t}^{\prime},\boldsymbol{\varepsilon}_{t}^{\prime\prime}$ are independent standard spatio-temporal white noises. Variability in $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ leads to non-stationarity (in space and time) of the conditional distribution ${\bf x}\,|(\boldsymbol{\theta},\boldsymbol{\Sigma})$, i.e., when $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ are simulated and fixed in the simulation of ${\bf x}_{t}$.
The reason for using this model was its ability to generate non-stationary random fields ${\bf x}_{t}$ without sacrificing linearity. The linearity of the forecast model allowed us to compare the filters with the unbeatable benchmark, the exact Kalman filter. This is normally not possible with nonlinear forecast models.
In the filters, the forecast model operator was ${\bf F}(\boldsymbol{\theta}_{t})$, and the forcing was zero. First, we generated the parameter fields $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ and computed the model error covariance matrices ${\bf Q}_{t}=\boldsymbol{\Sigma}_{t}\boldsymbol{\Sigma}_{t}^{\textsf{T}}$ (made available to the Kalman filter only). Then, $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ were fixed and we computed the truth ${\bf x}_{t}$ (as a solution to Eq. (21)) and observations ${\bf x}_{t}^{\rm obs}$ (as the noise-contaminated truth at some grid points and every analysis time instant $t$). After that, having fixed the truth and the observations, we ran the filters.
@@ 6.2 Filtering setup
The Doubly Stochastic Advection-Diffusion-Decay Model was defined on a 120-point regular grid on the circle (so that $\ell_{\rm max}=60$). Otherwise, its setup coincided with that reported in (Tsyrulnikov and Rakitko, 2019).
The ensemble size was between 5 and 160 with the default 10. The number of bandpass filters and the setup of the neural network were the same as in the above experiments on the sphere. 5000 consecutive assimilation cycles were used to validate the filters. ${\bf B}_{\rm mean}$ was computed by averaging Kalman filter’s ${\bf B}$ over 100,000 cycles and, additionally, by averaging in space. The latter was done by averaging entries of ${\bf B}$ over its main diagonal (getting the mean variance) and each cyclic super-diagonal (getting the whole set of mean covariances). “Synthetic” observations were located at every 10th model grid point. The observation error variance was specified to roughly match the mean Kalman-filter forecast error variance.
In EnKF and Hybrid-B filter, multiplicative covariance inflation (e.g., Houtekamer and Zhang (2016))) was used. Ensemble perturbations were multiplied by a factor slightly greater than one tuned to optimize the filter performance. The multiplier 1.02 was used in all experiments. There was no covariance inflation in LSEF (as we found that it deteriorated its performance, not shown).
The performance of the LSEF analysis reported in section 5 was so impressive that we did no tuning (other than choosing neural network’s hyperparameters) there. We recall that the reason for that was the use of the true prior model for $\sigma_{\ell}(x)$ (that is the true hyperprior model) in training the neural network. Here, as we had no true hyperprior model, we did some tuning. First, we found it beneficial to increase the number of bandpass filters from 6 to 8. Second, we specified the same width $\Delta_{H_{j}}=\Delta_{H}$ of the spectral transfer function (see Eq. (19)) for all filters. Third, we retrained the neural network for each ensemble size. Fourth, we increased $\Delta_{H}$ for larger ensemble sizes: $\Delta_{H}=5$ for $n_{\rm e}\leq 40$ and $\Delta_{H}=10$ for $n_{\rm e}>40$ (we recall that in I, we found that it is worth broadening the spectral transfer functions of the bandpass filters with the growing ensemble size). Fifth, we specified the (hyper)prior sample of the spectral functions $\sigma_{\ell}(x)$ in a different, ad-hoc way compared to section 5, see the next subsection.
@@ 6.3 Training sample
In LSEF, the forecast error field does not obey the Locally Stationary Convolution Model, its imposition in the LSEF analysis is an approximation. Therefore, in the course of filtering, we had no access to true local spectra (as they do not exist in this setting) and thus could not immediately apply the technique described in section 5.3 to train the neural network.
To build the training sample, we initially tried to manually set up a circular analogue of the Parametric Locally Stationary Convolution Model (described in J) in such a way that the spatial non-stationarity of the pseudo-random fields generated by it resembled those of the forecast-error fields we obtained in the course of Kalman filtering. In that way, we were able to generate the training sample following section 5.3. However, the filtering results were mediocre (not shown). Another option we tested was to fit, having the EnKF forecast ensemble at every assimilation cycle, the Locally Stationary Convolution Model. The fitted model provided us with the estimated local spectra, which were then used in the training sample as the “truth”. The respective sample band variances were computed from a new ensemble generated by the fitted model. This approach did yield an improvement, but the following technique was simpler and more successful, so it was used in the experiments presented below.
During an EnKF run, at each assimilation cycle $t=1,2,\dots,100,000$, we performed spatial averaging of the forecast-ensemble sample covariance matrices ${\bf S}_{t}$ (in the same way ${\bf B}_{\rm mean}$ was computed in section 6.2). The resulting spatially averaged covariance functions $b_{t}(x)$ (where $x=0,1,\dots,n_{\rm x}/2$ labels the distance measured in mesh sizes) were stored. Then, offline, for each covariance function $b_{t}(x)$ we computed its spectrum (Fourier transform) ${\bf f}_{t}=(f_{t}(0),f_{t}(1),\dots,f_{t}(\ell_{\rm max}))$. Knowing ${\bf f}_{t}$, we computed, for each $t$, an ensemble of pseudo-random realizations of the respective (stationary) random process, subjected them to the spatial bandpass filters (section 3.1), and finally computed the vectors ${\bf d}_{t}$ of sample band variances (section 3.2). The resulting training sample consisted of 100,000 pairs $({\bf d}_{t},\sqrt{{\bf f}_{t}})$. The above spatial averaging of the EnKF sample covariances reduced the diversity in the training sample. To compensate for that effect, we set up the Doubly Stochastic Advection-Diffusion-Decay Model in the strongly non-stationary regime (Tsyrulnikov and Rakitko, 2019, Table 3) when running EnKF to generate the training sample444 In a real-world application, say, in meteorology, the global averaging of EnKF sample covariances is to be replaced by local averaging to approximate a sample from the true hyperprior distribution. .
@@ 6.4 Results
The performance of each of the five filters was measured using the RMSE (with respect to the truth) of their control forecasts ${\bf x}^{\rm f}_{t}$. The RMSE was computed by averaging over the spatial grid, over 5000 assimilation cycles, and over 10 replicates of the parameter fields $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ of the model of truth. As in section 5.4, we show here the analysis performance score, see Eq. (20), where ${\rm RMSE}_{\rm TrueB}$ is the RMSE of the Kalman filter’s forecasts.
@@
Figure 6 shows that LSEF was better than the competing EnKF, Mean-B, and Hybrid-B filters for small to medium ensemble sizes. Even for large $M$, LSEF was superior to the competing filters. Comparing Fig. 6 with Fig. 2 demonstrates that the great advantage of LSEF observed when the truth was generated by the Parametric Locally Stationary Convolution Model (Fig. 2) was more modest in a more realistic situation here (Fig. 6). Nevertheless, the superiority of LSEF was significant for all ensemble sizes and substantial for $M\leq 40$. Note also that for a system with 120 degrees of freedom, the ensemble of size 160 is very large, a situation we never have in high-dimensional systems.
@@
Figure 7: Filters performance scores (the lower the better, =0 best possible), 90% bootstrap confidence bands (shaded). Dependence on the non-stationarity regime.
It is also worth noting that with each ensemble-based filter examined in our experiments, the spread (standard deviation) of the forecast ensemble did match the forecast RMSE pretty well. The difference was within 3% of the RMSE.
@@ 7 Discussion and conclusions
In this research, we have proposed the Locally Stationary Convolution Model and used it to regularize the observation-update (analysis) step of the ensemble Kalman filter.
@@
The spatial model is formulated (on the sphere and the circle) for the underlying forecast-error random field rather than for its spatial covariances, which guarantees that the implied covariance matrices are positive definite. The model is non-stationary in space, which can be essential for real-world (e.g., geoscientific) applications, and non-parametric, which has the potential of producing more diverse shapes of spatial correlations compared to those produced by parametric models. The spatial model is formulated as a process convolution model. The spatially varying convolution kernel $u(x,\rho)$ is constrained to reduce the effective number of degrees of freedom and regularize the estimation task.
@@
One essential regularizing constraint is local stationarity. It means that the random field (process) in question can be approximated by a stationary process locally in space. The non-trivial local stationarity takes place whenever the typical length scale of the process is significantly less than a non-stationarity length scale, the scale on which the local structure of the process varies in space. In other words, the convolution kernel $u(x,\rho)$ changes rapidly with the distance $\rho$ but slowly with the location $x$.
@@
The model allows for a definition of (spatially varying) local spectrum $f_{\ell}(x)=\sigma_{\ell}^{2}(x)$. Specifically, $\sigma_{\ell}(x)$ is the spectral transform of $u(x,\rho)$ with respect to $\rho$. Another fundamental constraint imposed on the spatial model is formulated in spectral space as smoothness of $\sigma_{\ell}(x)$ with respect to $\ell$. This constraint means that there is no need for high spectral resolution in estimating $\sigma_{\ell}(x)$. The relatively low spectral resolution allows for a higher resolution in physical space, thus permitting to capture spatial non-stationarity in the model estimator. Besides, smooth local spectra imply rapidly decaying spatial correlations. As a result, spurious long distance correlations are suppressed, which is important in data assimilation applications.
@@
In general, using a spatial model to regularize the analysis step in a high-dimensional ensemble filter is a way to incorporate the available problem-specific spatial information into the analysis. Theoretically, the constraints are part of the (hyper)prior model for $\sigma_{\ell}(x)$ used by the spatial-model estimator. The list of constraints can be modified or extended. For example, we may know the typical degree of roughness/smoothness of the analyzed field (determined by the behavior of spatial correlations at small distances). With our spatial model, this knowledge can be encoded in a constraint on the rate of decay of the local spectra at large wavenumbers.
@@
Aiming at high-dimensional applications, we built a computationally efficient model estimator. Relying on the local stationarity and smooth-local-spectra constraints, the estimator employs spatial bandpass filters to extract aggregated local spectra from the input ensemble of random field realizations. The aggregated local spectra are disaggregated (deconvolved) by a neural Bayes estimator. Theoretically, we have shown that the bandpass filters can be specified in such a way that the resulting model estimator is consistent.
@@
Due to the regularizing constraints, the estimator produces much less noise in the resulting spatial covariances compared with traditional localized (tapered) sample covariances. However, it does so at the expense of introducing a systematic distortion in the covariances/spectra—because the imposed constraints need not be respected in real-world spatial processes. The main hypothesis underlying this research is that there are filtering problems in which the reduced noise in the spatial covariances (i.e., the reduced random uncertainty) has a greater positive impact on the filter’s accuracy than the negative effect of the induced systematic uncertainty in the filter’s prior distribution. Our numerical experiments suggest that this hypothesis is plausible.
@@
We postulated that in the course of sequential ensemble filtering, the forecast error field obeys the Locally Stationary Convolution Model. As a result, we obtained a new sequential filter termed the Locally Stationary convolutional Ensemble Filter (LSEF). Testing the filter with a non-stationary spatio-temporal model of truth yielded encouraging results. It is worth noting that ad-hoc regularization devices like localization and covariance inflation, which are, normally, used (and employed in our experiments) in EnKF and hybrid filters were absent in LSEF. This can be important in practice since ad-hoc devices require extensive (and expensive) tuning. Another interesting result is that the advantage of LSEF was observed not only under weak non-stationarity, as intended, but in strongly non-stationary regimes as well.
@@
The tests of our spatial-model-based covariance regularization technique were conducted with linear observation models on the sphere and linear forecast and observation models on the circle. We believe that linearity of these tests is not a significant restriction, expecting that the technique will be useful whenever an EnKF-like filter is useful, that is, with non-linear models as well.
@@
Using a neural Bayes estimator to recover local spectra from the outputs of the spatial bandpass filters proved to be successful even in the situation when true (target) local spectra were not available to train the net (in sequential filtering). In case when the target local spectra were available (in our experiments with static analyses on the sphere) and they were used to build the training sample, the advantage of the neural network was dramatic. Our results also suggest that the neural net was more successful in extracting “climatological” information from the training sample and combining it with the information provided by the forecast ensemble than the widely-used linear combination of static and localized ensemble covariance matrices.
Applying the developed regularization technique to real-world high-dimensional problems will require a multi-grid/multi-scale version of the algorithm. Work on it is underway. Its non-ensemble version (with static spatial covariances) has been successfully tested in an operational meteorological data assimilation system (Tsyrulnikov et al., 2019). Adding the ensemble component should be affordable because the spatial model estimator consists of computationally cheap spatial filters followed by a grid point-by-grid point application of a small neural network. Further extensions of the spatial model may include directional anisotropy and, possibly, non-Gaussianity. In applications to non-global domains, spectral bandpass filtering can be replaced with spatial wavelet filtering.

### v5

@@ 6.6 Effects of covariance regularization on spatial covariances
Figure 6(left) shows that the sample (EnKF’s) variances are much noisier than the model variances. This illustrates the fact that the EnKF’s covariance localization cannot improve sample variances, which are quite uncertain for small ensemble sizes. Less noise in the model (LSEF) variances is due to the regularizing effect of the constrained spatial model. In particular, the local stationarity constraint implicitly allows for the use of neighboring covariances in the estimation. The specific mechanism for that is provided by the LSEF’s bandpass filters.
The right panel of Fig. 6 shows that, similar to the left panel, the model correlations contain much less noise than their localized sample counterparts. Another feature seen in this figure is that, in contrast to model correlations, localized sample correlations suffer from a negative bias in their length scale (see section 1.1, item 1). Furthermore, Fig. 6(right) illustrates that fact that the LSEF analysis produces almost noise-free large-distance background-error correlations without any ad-hoc localization (tapering) device. This emphasizes the usefulness of the smooth-local-spectrum constraint (item (4) in section 3.8) and the ability of the neural Bayes estimator to ensure this constraint is met. One can also notice in Fig. 6(right) that the regularizing constraint (1) (section 3.8), which requires the kernel to be locally isotropic, does not lead to isotropic correlations (note the asymmetry in both the true and model correlations).
Averaged over 300 realizations of the truth and the ensemble, the mean absolute error in the model (LSEF) variances was about 1.5 times less than the respective error in the sample variances. The mean absolute error in the model (LSEF) spatial correlations was about half of the error in the localized sample correlations. So, both LSEF’s variances and correlations were substantially more accurate than their localized sample counterparts. This explains the significant advantage of the LSEF analysis over the EnKF analysis in the experiments presented above in this section.
@@ 6.7 Conclusions on the experiments with static analyses
When the truth was generated by a (parametric) Locally Stationary Convolution Model, the proposed analysis technique performed very well compared with the competing techniques. The question remains how the proposed approach performs in a more realistic situation when background errors do not obey the Locally Stationary Convolution Model and the true local spectra are not just unavailable but they do not exist. To answer this question we turn to sequential data assimilation.
@@ 7.1 Model of truth and forecast model
${\bf x}_{t}={\bf F}(\boldsymbol{\theta}_{t}){\bf x}_{t-1}+\boldsymbol{\Sigma}_{t}\boldsymbol{\eta}_{t},$ (35)
where ${\bf x}_{t}$ is the spatial field defined on a regular grid on the unit circle (a vector), $t$ is the time index, ${\bf F}(\boldsymbol{\theta}_{t})$ is the forecast model operator (a matrix), $\boldsymbol{\theta}_{t}$ and $\boldsymbol{\Sigma}_{t}$ are the spatial ‘parameter’ fields (vectors) defined below in Eqs. (36) and (37), $\boldsymbol{\eta}_{t}$ is the standard spatio-temporal white noise, and multiplication in the second term on the right-hand side is component-wise. Double stochasticity means that both the forcing $\boldsymbol{\eta}_{t}$ and the parameters $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ are random. $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ are postulated to be stationary (in space and time) random fields satisfying the same model Eq. (35) but with constant non-random coefficients $\theta_{\theta},\theta_{\Sigma},\Sigma_{\theta},\Sigma_{\Sigma}$:
$\boldsymbol{\theta}_{t}={\bf F}(\theta_{\theta})\boldsymbol{\theta}_{t-1}+\Sigma_{\theta}\boldsymbol{\eta}_{t}^{\prime},$ (36)
$\boldsymbol{\Sigma}_{t}={\bf F}(\theta_{\Sigma})\boldsymbol{\Sigma}_{t-1}+\Sigma_{\Sigma}\boldsymbol{\eta}_{t}^{\prime},$ (37)
where $\boldsymbol{\eta}_{t}^{\prime},\boldsymbol{\eta}_{t}^{\prime}$ are independent standard spatio-temporal white noises. Variability in $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ leads to non-stationarity (in space and time) of the conditional distribution ${\bf x}\,|(\boldsymbol{\theta},\boldsymbol{\Sigma})$, i.e., when $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ are simulated and then fixed in the simulation of ${\bf x}_{t}$.
The reason for using this model was its ability to generate non-stationary random fields ${\bf x}_{t}$ without sacrificing linearity. Linearity of the forecast model allowed us to compare the filters with the unbeatable benchmark, the exact Kalman filter. This is normally not possible with nonlinear forecast models.
In the filters, the forecast model operator was ${\bf F}(\boldsymbol{\theta}_{t})$ and the forcing was zero. First, we generated the parameter fields $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ and computed the model error covariance matrices ${\bf Q}_{t}=\boldsymbol{\Sigma}_{t}\boldsymbol{\Sigma}_{t}^{\textsf{T}}$ (made available to the Kalman filter only). Then, $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ were fixed and we computed the truth ${\bf x}_{t}$ (as a solution to Eq. (35)) and observations ${\bf x}_{t}^{\rm obs}$ (as the noise-contaminated truth at some grid points and every analysis time instant $t$). After that, having fixed the truth and the observations, we ran the filters.
@@ 7.2 Filtering setup
The Doubly Stochastic Advection-Diffusion-Decay Model was defined on a 120-point regular grid on the circle (so that $\ell_{\rm max}=60$). Otherwise, its setup coincided with that reported in Tsyrulnikov and Rakitko (2019).
The ensemble size was between 5 and 160 with the default 10. The number of bandpass filters and the setup of the neural network were the same as in the above experiments on the sphere. 5000 consecutive assimilation cycles were used to validate the filters. ${\bf B}_{\rm mean}$ was computed by averaging Kalman filter’s ${\bf B}$ over 100,000 cycles and, additionally, by averaging in space. The latter was done by averaging entries of ${\bf B}$ over its main diagonal (getting the mean variance) and each cyclic super-diagonal (getting the whole set of mean covariances). ‘Synthetic’ observations were located at every 10th model grid point. The observation error variance was specified to roughly match the mean Kalman-filter forecast error variance (this resembles the relationship between observation and forecast errors in meteorological data assimilation, our application area).
In the EnKF and Hybrid-B filters, multiplicative covariance inflation (e.g. Houtekamer and Zhang, 2016) was used and tuned to optimize their performance. Ensemble perturbations were multiplied by a factor of 1.02. There was no covariance inflation in LSEF (as we found that it deteriorated its performance, not shown).
The performance of the LSEF analysis reported in section 6 was so impressive that we did no tuning there (other than choosing the neural network’s hyperparameters). We recall that the reason for that was the setting in which (i) the prior uncertainty was, by construction, governed by the Locally Stationary Convolution Model and (2) the hyperprior of the model parameters was also available to train the neural network. Here, as neither of these two conditions was met, some tuning was done. First, we found it beneficial to increase the number of bandpass filters from 6 to 8. Second, we specified ${q_{\rm shape}}=2$ and the same width $\Delta_{{\cal H}_{j}}=\Delta_{{\cal H}}$ of the spectral transfer function (see Eq. (33)) for all bandpass filters. Third, we retrained the neural network for each ensemble size. Fourth, we increased $\Delta_{{\cal H}}$ for larger ensemble sizes: $\Delta_{{\cal H}}=5$ for $n_{\rm e}\leq 40$ and $\Delta_{{\cal H}}=10$ for $n_{\rm e}>40$ (following the finding in G that it is worth broadening the spectral transfer functions of the bandpass filters with the growing ensemble size, see also section 4.4). Fifth, we specified the (hyper)prior sample of the spectral functions $\sigma_{\ell}(x)$ in a different, ad-hoc way compared to section 6, see the next subsection.
@@ 7.3 Training sample
In the course of sequential filtering with LSEF, the forecast error field does not have to obey the Locally Stationary Convolution Model, its imposition in the LSEF analysis is an approximation. Therefore we had no access to true local spectra here (as they do not exist in this setting) and thus could not immediately apply the technique described in section 6.4 to train the neural network.
To build the training sample, we initially tried to manually set up a circular analogue of the Parametric Locally Stationary Convolution Model (described in H) in such a way that the spatial non-stationarity of the pseudo-random fields generated by it resembled those of the forecast-error fields we obtained in the course of Kalman filtering. In that way, we were able to generate the training sample following section 6.4. However, the filtering results were mediocre (not shown). Another option we tested was to fit, having the EnKF forecast ensemble at every assimilation cycle, the Locally Stationary Convolution Model. The fitted model provided us with the estimated local spectra, which were then used in the training sample as the ‘truth’. The respective sample band variances were computed from a new ensemble generated by the fitted model. This approach did yield an improvement, but the following technique was simpler and somewhat more successful, so it was used in the experiments presented below.
During an EnKF run, at each assimilation cycle $t=1,2,\dots,10^{5}$, we performed spatial averaging of the forecast-ensemble sample covariance matrices ${\bf S}_{t}$ (in the same way the spatial averaging was done in the computation of ${\bf B}_{\rm mean}$, see section 7.2). The resulting spatially averaged covariance functions $b_{t}(x)$ (where $x=0,1,\dots,n_{\rm x}/2$ labels the distance measured in mesh sizes) were stored. Then, offline, for each covariance function $b_{t}(x)$ we computed its spectrum (Fourier transform) ${\bf f}_{t}=(f_{t}(0),f_{t}(1),\dots,f_{t}(\ell_{\rm max}))$. Knowing ${\bf f}_{t}$, we computed, for each $t$, an ensemble of pseudo-random realizations of the respective (stationary) random process, subjected them to the spatial bandpass filters (section 4.1), and finally computed the vectors ${\bf d}_{t}$ of sample band variances (section 4.3). The resulting training sample consisted of 100,000 pairs $({\bf d}_{t},\sqrt{{\bf f}_{t}})$. The above spatial averaging of the EnKF sample covariances reduced the diversity in the training sample. To compensate for that effect, we set up the Doubly Stochastic Advection-Diffusion-Decay Model in the strongly non-stationary regime (Tsyrulnikov and Rakitko, 2019, Table 3) when running EnKF to generate the training sample.
@@ 7.4 Results
The performance of each of the five filters was measured using the RMSE (with respect to the truth) of their control forecasts ${\bf x}^{\rm f}_{t}$. The RMSE was computed by averaging over the spatial grid, over 5000 assimilation cycles, and over 10 replicates of the parameter fields $\boldsymbol{\theta}_{t},\boldsymbol{\Sigma}_{t}$ of the model of truth (outlined in section 7.1). As in section 6.5, we show here the analysis performance score, Eq. (34), where ${\rm RMSE}_{\rm TrueB}$ is the RMSE of Kalman filter forecasts.
@@
Figure 7 shows that LSEF was significantly better than the competing EnKF, Mean-B, and Hybrid-B filters. Comparing Fig. 7 with Fig. 3 demonstrates that the great advantage of LSEF observed when the truth was generated by the Parametric Locally Stationary Convolution Model (Fig. 3) was more modest in a more realistic situation here (Fig. 7). Nevertheless, the LSEF outperformed the three competing approximate filters even for very large ensemble sizes. Note that for a system with 120 degrees of freedom, the ensemble of size 160 is indeed very large, a situation we never have in high-dimensional systems. For $K\leq 40$, the superiority of LSEF was substantial.
@@
Figure 8: Filters performance scores (the lower the better, =0 best possible), 90% bootstrap confidence bands (shaded). Dependence on the non-stationarity regime (0 is stationary, 1 is weakly non-stationary, 2 is default, 3 is highly non-stationary).
It is also worth mentioning that with each ensemble-based filter examined in our experiments, the spread (standard deviation) of the forecast ensemble did match the forecast RMSE pretty well. The difference was within 3% of the RMSE.
@@ 8 Discussion and conclusions
This research aims to enhance the ensemble Kalman filter by introducing a spatial model at the analysis (observation update) step. The idea is to abandon the use of sample covariances, which are notoriously noisy in high-dimensional problems. Instead, we propose to build a non-parametric, non-stationary spatial model, constrain it to reduce its number of degrees of freedom, and estimate it directly from the prior ensemble. The estimated model is then used to specify prior covariances in the analysis.
@@
The model is formulated (on the sphere and the circle, but can be extended to other domains) for the underlying forecast-error random field rather than for its spatial covariances, which guarantees that the implied covariance matrices are positive definite. The model is non-stationary in space, which can be essential for real-world (e.g., geoscientific) applications, and non-parametric, which has the potential of producing more diverse shapes of spatial correlations compared to those produced by parametric models. The spatial model is formulated as a Gaussian process convolution model. The spatially varying convolution kernel $u(x,\rho)$ is constrained to reduce the effective number of degrees of freedom and regularize the estimation task.
@@
One fundamental regularizing constraint is local stationarity, hence the name, the Locally Stationary Convolution Model. Local stationarity means that the random field (process) in question can be approximated by a stationary process locally in space. Local stationarity takes place whenever the typical length scale of the process is significantly less than a non-stationarity length scale, the scale on which the local structure of the process varies in space. In other words, whenever the convolution kernel $u(x,\rho)$ changes slowly with location ($x$) compared with its change as a function of distance ($\rho$). The stationary process-convolution model is a special case of our non-stationary model.
@@
Another fundamental constraint imposed on the spatial model is formulated in spectral space for spectral functions $\sigma_{\ell}(x)$ defined to be the spectral transform of $u(x,\rho)$ with respect to $\rho$. We require $\sigma_{\ell}(x)$ (or the local spectra $f_{\ell}(x)=\sigma^{2}_{\ell}(x)$) to be smooth functions of the wavenumber $\ell$. This constraint means that there is no need for high spectral resolution in estimating the spectral functions $\sigma_{\ell}(x)$. The relatively low spectral resolution allows for a higher resolution in physical space, thus permitting to capture spatial non-stationarity in the model estimator. Besides, smooth local spectra imply rapidly decaying spatial correlations. As a result, spurious long distance correlations are suppressed (without additional ad-hoc devices like localization or tapering), which is important in data assimilation applications.
@@
The constraints are part of a (hyper)prior model for the model parameters. The list of constraints can be modified or extended. For example, we may know the typical degree of roughness/smoothness of the analyzed field (determined by the behavior of spatial correlations at small distances). With our spatial model, this knowledge can be encoded in a constraint on the rate of decay of the spectral functions $\sigma_{\ell}(x)$ at large wavenumbers.
In general, using a spatial model to regularize the analysis step in a high-dimensional ensemble filter is a way to incorporate problem-specific spatial information into the analysis. This information may be available or can be acquired by a dedicated effort.
@@
Aiming at high-dimensional applications, we built a computationally efficient model estimator. Relying on the local stationarity and smooth-local-spectra constraints, the estimator employs spatial bandpass filters to extract aggregated local spectra from the input ensemble of random field realizations. The aggregated local spectra are then disaggregated (deconvolved) from sample variances of the filtered fields. With a linear disaggregation technique, we prove that the bandpass filters can be specified in such a way that the resulting model estimator is consistent. A more accurate nonlinear disaggregation method based on the use of a neural network is developed.
@@
Due to the regularizing constraints, the estimator produces much less noise in the resulting spatial covariances compared with the traditional localized (tapered) sample covariances. However, it does so at the expense of introducing a systematic distortion in the covariances/spectra—because the imposed constraints need not be respected in real-world spatial processes. The main hypothesis underlying this research is that there are filtering problems in which the reduction of noise in spatial covariances (i.e., the reduced random uncertainty) has a greater positive impact on the filter’s accuracy than the negative effect of the induced systematic uncertainty in the filter’s prior distribution. Our numerical experiments suggest that this hypothesis is plausible.
@@
We postulated that in the course of sequential ensemble filtering, the prior uncertainty is governed by the Locally Stationary Convolution Model estimated online from the prior ensemble and a hyperprior distribution. This allowed us to optimize the use of both the ensemble and the forecast-error ‘climatology’. As a result, we obtained a new sequential filter termed the Locally Stationary convolutional Ensemble Filter (LSEF). Testing the filter with a non-stationary spatio-temporal model of truth yielded encouraging results. It is worth noting that ad-hoc regularization devices like localization and covariance inflation, which are, normally, used in EnKF and hybrid filters were absent in LSEF. This can be important in practice since ad-hoc devices require extensive (and expensive) tuning. Another interesting result is that the advantage of LSEF was observed not only under weak non-stationarity, as intended, but in strongly non-stationary regimes as well.
@@
Our spatial-model-based covariance regularization technique was tested with linear observation models on the sphere and linear forecast and observation models on the circle. We believe that linearity of these tests is not a significant restriction, expecting that the technique will be useful whenever an EnKF-like filter is useful, that is, with non-linear models as well.
@@
Using a neural Bayes estimator to recover local spectra from the outputs of the spatial bandpass filters proved to be successful even in the situation when true hyperprior distribution was not available to train the neural net (in sequential filtering). In case when the true hyperprior was available (in our experiments with static analyses on the sphere), the advantage of the neural-Bayes-estimator based analysis was dramatic. Our results also suggest that placing a hyperprior on the spatial model’s parameters leads to better extraction of ‘climatological’ information from the training sample than the mixing of ensemble prior covariances with the time-mean prior covariance matrix (the popular technique reviewed in section 1.1, see item 2). One reason for this is that the hyperprior distribution is more informative than just the mean covariance matrix. Another reason is that the expressivity of a neural network is much greater than that of a linear combination with scalar weights.
Applying the developed regularization technique to real-world high-dimensional problems will require a multigrid/multi-scale version of the algorithm. Work on it is underway. The key idea is to approximate a given convolution operator using sparse operators defined on a hierarchy of spatial grids. The computational cost will be drastically reduced by representing large-distance spatial interactions on coarse grids.
Adapting our model to fields with strong non-stationarities that occur at fixed known locations (like ocean coastlines) can also be subject of future research. Further extensions of the spatial model may include directional anisotropy, and, possibly, non-Gaussianity. In applications to non-global domains, spectral bandpass filtering can be replaced with spatial wavelet filtering.
Last but not least, a promising avenue for further research is to extend the approach proposed in this paper to deal with imperfect prior ensembles. This can be especially important if the ensemble is generated by a fast data-driven forecast model. Even though such an ensemble can be large, it can suffer from probability biases. The role of the spatial model will then be to correct for those biases.